How much of a pure 100% acid must be added to 2 liters of a 10% acid solution in order to produce a 20% solution? How much 20% solution will there be?
Answer by mananth(16946)

------------ percent ---------------- Amount
Acid type I 10 ---------------- 2 liters
Pure acid 100 ---------------- x liters
Total 20 ---------------- 2 + x liters

10*2+100x=20(2 +x)
20+100 x=40+20 x
100x-20 x =-20+ 40
80 x = 20
/ 80
x = 0.25 of liters Pure acid
0.25 of liters Acid type I
m.ananth@hotmail.ca